OpenStudy (anonymous):

Let f(x) = x^(4/3) - 7 x^(-8/3) f '(x) = A xB + C xD, where I think they want me to find a,b,c & d

OpenStudy (anonymous):

I think I use the power rule

OpenStudy (turingtest):

exactly, and what dd you get?

OpenStudy (anonymous):

f(x)= rx^r-1

OpenStudy (anonymous):

and I get ...

OpenStudy (anonymous):

4/3x^1/3- -56/3x^-3/2 is that right?

OpenStudy (turingtest):

first term is right, but not the second what is\[-\frac83-1\]?

OpenStudy (anonymous):

-7/3

OpenStudy (turingtest):

\[-\frac83-1=-\frac83-\frac33=-\frac{11}3\]which is your exponent on the second term

OpenStudy (anonymous):

so 4/3x^1/3- -56x^-11/3

OpenStudy (turingtest):

negative times negative is positive we never write -- and you forgot to divide 56/3 in your final answer

OpenStudy (anonymous):

oh yeah thanks didnt see that

OpenStudy (anonymous):

wait something is wrong 56x^11/3

OpenStudy (anonymous):

i mean 56x^-11/3

OpenStudy (turingtest):

(56/3)x^(-11/3)

OpenStudy (anonymous):

man I need to be careful

OpenStudy (turingtest):

-7x^(-8/3) -7(-8/3)x^(-8/3-1) 56/3x^(-11/3) yeah, algebra can get you if you're not paying attention :P